Droz's Hiking Gear, Inc. has found that its common equity capital shares have a beta equal to 1.5 while the risk-free return is 8 percent and the expected return on the market is 14 percent. It has 7-year semiannual maturity bonds outstanding with a price of $767.03 that have a coupon rate of 7 percent. The firm is financed with $120,000,000 of common shares (market value) and $80,000,000 of debt. What is the after-tax weighted average cost of capital for Droz's, if it is subject to a 35 percent marginal tax rate? Round your final percentage answer to two decimal places.

Show in Excel pls thanks!
Answer #1

Equity Beta (be) = 1.5

Risk Free Rate (Rf) = 8%

Expected Market Return (Rm) = 14%

Using Capital Asset Pricing Model

Ke = Rf + be * (Rm - Rf)

Ke = 8% + 1.5 * (14% - 8%) = 17%

Now Calculation of Cost of Debt (Kd) [ Since not given in the question, lets assume face value of bond = $1,000)

Interest Paid = $1,000 * 7% = $70

Market Price of Bond = $767.03

Kd = $70 / $767.07 (Pre Tax)

Kd = 9.13% (Pre Tax)

Kd (Post Tax) = 9.13% * (1 - 0.35) = 5.93%

Now Calculation of Weighted Average Cost of Capital (After Tax)

WACC =(Kd * Wd) + (Ke + Wd)

Wd = 80,000,000/(80,000,000 + 120,000,000) = 0.4

We = 1 - Wd = 0.6

WACC = 5.93% * 0.4 + 17% * 0.6 = 12.57%